OpenStudy (anonymous):

if im given a table with data, how do i find the line of best fit? theres no graph btw.

OpenStudy (abb0t):

plot the data on a graph paper.

OpenStudy (abb0t):

draw a line in such a way that the line touches most points.

OpenStudy (anonymous):

i cant use graph paper D':

OpenStudy (abb0t):



OpenStudy (anonymous):

thank you, but i cant use graph paper on my test tommorow....

OpenStudy (anonymous):

Is it a stats question?

OpenStudy (anonymous):

Temp. (degrees) 10 50 40 30 Rainfall (mm) 80 90 100 90 this is the graph im looking at

OpenStudy (anonymous):

usually we are given a graph but it says to calculate the equation for the line of best fit...

OpenStudy (tkhunny):

You must define a Fit Measure, otherwise "best" doesn't mean anything. Are you working on a section on Linear Regression or Least Squares?

OpenStudy (amistre64):

a linear best fit (linear regression models) take the averageX and averageY to define the "point" of the line with then the slope of the line is a bit more calcuation extensive but follows the basic setup:\[\frac{n\sum xy-\sum x\sum y}{n\sum xx-\sum x\sum x}\]

OpenStudy (amistre64):

xx: 100 2500 1600 900 x: Temp. (degrees) 10 50 40 30; sum x = 130 y: Rainfall (mm) 80 90 100 90; sum y = 360 xy: 800 4500 4000 2700 and since there are 4 points, n=4

OpenStudy (anonymous):

thanks so much :)

OpenStudy (amistre64):

youre welcome